1. y'-2y=y^2 ;y=3 when x=0
2. (y')^2=(1-y^2)/(1-x^2); y=1/2 when x=1
3. x(e^y)dy+((x^2)+1dx)/y=0
1
Expert's answer
2020-10-12T18:58:33-0400

1. Rewrite the equation:


"y' = y^2 +2y""\frac{dy}{dx}=y^2+2y""\frac{dy}{y^2+2y}=dx""\frac{dy}{(y^2+2y+1)-1}=dx""\frac{dy}{(y+1)^2-1}=dx"


Integrate both parts of the equation:


"\int \frac{dy}{(y+1)^2-1}= \int dx""\frac{1}{2} ln |\frac{y}{y+2}| =x+C""ln |\frac{y}{y+2}| =2x+C""\frac{y}{y+2}=e^{2x+C}""y=e^{2x+C}y+2e^{2x+C}""y(1-2e^{2x+C})=2e^{2x+C}""y=\frac{2e^{2x+C}}{1-2e^{2x+C}}"

Substitute y=3, x=0:


"3 = \frac{2e^C}{1-2e^C}""2e^C =3-6e^C""8e^C =3""e^C =\frac{3}{8}"

Hence


"y= \frac{2 \cdot \frac{3}{8}e^{2x}}{1-2 \cdot \frac{3}{8}e^{2x}}""y= \frac{ \frac{3}{4}e^{2x}}{1-\frac{3}{4}e^{2x}}""y= \frac{ 3e^{2x}}{4-3e^{2x}}"

2. Rewrite the equation:


"(\frac{dy}{dx})^2 =\frac{1-y^2}{1-x^2}""\frac{dy}{dx}=\frac{\sqrt{1-y^2}}{\sqrt{1-x^2}}""\frac{dy}{\sqrt{1-y^2}}=\frac{dx}{\sqrt{1-x^2}}"

Integrate both parts of the equation:


"\int \frac{dy}{\sqrt{1-y^2}}=\int \frac{dx}{\sqrt{1-x^2}}""arcsin y = arcsin x +C"

Since "y = \frac{1}{2}, x = 1," then


"arcsin \frac{1}{2} = arcsin {1} +C""C = \frac{\pi}{6}-\frac{\pi}{2} =-\frac{\pi}{3}"

Hence


"arcsin y = arcsin x -\frac{\pi}{3}""y = sin ( arcsin x -\frac{\pi}{3})""y = x cos \frac{\pi}{3}-sin \frac{\pi}{3} cos (arcsin x)""y = \frac{x}{2}-\frac{\sqrt {3}}{2}\cdot \sqrt {1-x^2}""y = \frac{1}{2}(x- \sqrt {3-3x^2})"

3. Rewrite the equation:


"x e^y dy = -\frac{x^2+1}{y}dx""xy e^y dy = -(x^2+1)dx""y e^y dy = -\frac{x^2+1}{x}dx""y e^y dy = -(x+\frac{1}{x})dx"

Integrate both parts of the equation:


"\int{y e^y dy} = -\int (x+\frac{1}{x})dx""y e^y -e^y = -\frac{x^2}{2}+ln x+C""e^y(y-1)= -\frac{x^2}{2}+ln x+C"